OpenStudy (anonymous):

Write an equation of the line that is parallel to y = -2x + 6 and passes through (3,2) .

OpenStudy (anonymous):

parallel meaning that they have the same gradient gradient of unknown must be -2. and passes through the point (3,2) so use the basic linear equation y=mx+c m=gradient and c-intercept. substitute the gradient and the point into the general equation to get ur c value and then re-write the whole equation with m and c.

OpenStudy (anonymous):

tell me ur answer after u've got it, I will check it for u.

OpenStudy (anonymous):

Substitute your values from the question into this equation. \[y-y _{1}=m(x-x _{1})\] Where the point \[(x _{1}, y _{1})\] is the point \[(3, 2)\] and m is the gradient -2.